### v3

# Efficiently Computing Equilibria in Budget-Aggregation Games

Abstract
Budget aggregation deals with the social choice problem of distributing an exogenously given budget among a set of public projects, given agents’ preferences. Taking a game-theoretic perspective, we study budget-aggregation games where each agent has virtual decision power over some fraction of the budget. We investigate the structure and show efficient computability of Nash equilibria for various common preference models in this setting. In particular, we show that equilibria for Leontief utilities can be found in polynomial time, solving an open problem from Brandt et al. (2023), and give an explicit polynomial-time algorithm for computing equilibria for $\ell_{1}$ preferences.
1 Introduction
Participatory budgeting (Cabannes, 2004; Aziz and Shah, 2021) is a democratic process in which citizens are allowed to vote on how the funds of their city or state should be distributed. It is justified both by ethical arguments (people have a fundamental right to decide how their money is used) and by practical arguments (letting people decide on the budget strengthens cohesion in the community).
This approach can be taken to the extreme: instead of just letting citizens vote on which projects they approve (as is commonly done today), one can give each citizen their own (virtual) share of the budget (e.g. $1/n$ of the total budget, in case there are $n$ citizens with equal entitlements), and allow them to decide how to allocate that share among the potential projects. This can be an attractive approach, as it gives full power to the citizens and ensures some degree of fairness by splitting this power equally among them. Furthermore, each citizen is able to directly observe her influence on the outcome by means of the distribution of her individual budget.
However, this approach raises severe strategic problems. For example, if a citizen thinks that projects $x$ and $y$ should each receive $0.5$ of the total budget, but later observed that most other citizens allocated their shares to $x$, then she would have preferred to spend her entire budget on $y$, leaving her regretful since her budget was spent sub-optimally in hindsight. This bad user experience can be avoided if we have the possibility to coordinate individual budgets by advising each citizen in advance how to distribute their own budget such that for each citizen, the advised distribution is optimal, given that all other citizens follow the provided advice.
In other words, we would like to compute a Nash equilibrium (NE) of the game in which the players are citizens, and the set of strategies of each player is the set of possible distributions of their share of the budget. We call this game the budget-aggregation game. An NE of the budget-aggregation game hits a sweet spot between giving complete autonomy to the citizens to distribute their shares of the budget and ensuring that they are indeed satisfied with their own choice. More specifically, every agent receives at least their maximin share, a property studied in fair division (see, e.g., Brandt et al., 2016). In words, an agent receives the highest possible utility they can secure with their share of the budget under the least favorable allocation of the remaining shares.
Our contributions
We propose a new class of normal-form games, called budget-aggregation games, which model (equal-resource) public funding situations. We position these games within the broader framework of normal-form games and survey existing results on the computation of NEs. On top of that, our paper analyzes NEs for the most common utility functions that have been considered in participatory budgeting and related fields. We prove theorems on the structure of NEs and develop algorithms to efficiently compute them. We focus on the following utility models:
  • For linear utilities, we show that an NE can be computed by solving a linear program (Section 4).
  • For Leontief utilities, we show that an NE can be computed in polynomial time using a variant of the ellipsoid method, solving an open problem from Brandt et al. (2023) (Section 5).
  • For binary symmetric separable utilities, we explore strategic similarities to Leontief utilities to show that an NE can be found in polynomial time. (Section 6).
  • For $\ell_{1}$ disutilities, we again present an intuitive polynomial-time algorithm to compute an NE (Section 7).
Finally, we extend the basic model by allowing agents to have different weights. We show that, for most of the above utility models (namely linear, Leontief and $\ell_{1}$), an NE can still be computed in pseudo-polynomial time (Section 8).
2 Related Work
About $75$ years ago, two groundbreaking papers by John Nash (Nash, 1950, 1951) proved the existence of an NE in general normal-form games. Apart from its intrinsic value, the proof technique of applying fixed-point theorems, especially Kakutani’s fixed-point theorem (which Nash attributed to David Gale), paved the way for other equilibrium existence theorems like Debreu’s social equilibrium (Debreu, 1952) or general equilibrium in the Arrow-Debreu model (Arrow and Debreu, 1954; McKenzie, 1959).
An important generalization of Nash’s theorem involves the notion of a concave game (Rosen, 1965). A concave game is defined by, for each player $i$, a convex set $(S_{i})_{i\in N}$ that represents the possible strategies for $i$, and a utility function $u_{i}$ that maps each strategy profile to a real number; $u_{i}$ can be any function that is continuous in all strategies, and concave in the strategy of ii. [^1] The definition in Rosen (1965) is even more general and allows coupled constraints, that is, the set of strategies available to each agent may depend on the choices made by other agents. We do not need this generalization here. Nash’s game is a special case in which every $S_{i}$ is a standard simplex in some Euclidean space (representing the set of all lotteries over pure actions), and every $u_{i}$ is a linear combination of products of $n$ terms. Our budget-aggregation game is a special case in which every $S_{i}$ is (still) a simplex, but the utility functions are substantially different, due to the interpretation as a distribution of exogenously given budget rather than a lottery.
2.1 Computing an equilibrium is hard
Like most fixed-point-based proofs, the ones by Nash (1950, 1951) and Rosen (1965) do not yield efficient algorithms for finding an equilibrium. Moreover, there are 3-player games with integer payoffs with a unique (mixed) NE in which all probabilities are irrational numbers (Nash, 1951; Bilò and Mavronicolas, 2014), and 4-player games with a unique NE in which all probabilities are irradical numbers (cannot be expressed using integer roots) (Orzech and Rinard, 2025). This shows that no finite algorithm can compute an exact NE in general games.
Even computing an $\epsilon$-approximate NE probably cannot be done in polynomial time, as it is PPAD-hard even for two players. This was first proven for $\epsilon$ that shrinks exponentially in the number of actions (Chen and Deng, 2006), then for $\epsilon$ that shrinks polynomially (Chen et al., 2006), and later even for some constant $\epsilon>0$ (Rubinstein, 2015). Computing an equilibrium is PPAD-complete, also for some other classes of concave games (Papadimitriou et al., 2023). Many natural decision problems regarding the existence of a NE with certain properties are NP-hard for two-player games, and $\exists\mathbb{R}$-hard for multi-player games (Bilò and Mavronicolas, 2014).
However, there are other classes of games for which an NE can be computed efficiently. We survey some of them below.
2.2 Nash games
For the original game studied by Nash (where the strategy set of each agent is the set of lotteries over a finite set of actions), many practical algorithms for computing an NE were designed, such as the ones by Lemke and Howson (1964) and Porter et al. (2008), but without a polynomial time guarantee.
Polynomial-time algorithms for $\epsilon$-approximate NE (where no agent can increase her utility by more than $\epsilon$ by deviating, where the utilities are normalized to $[0,1]$) are known for $\epsilon\approx 0.34$ for two players (Tsaknakis and Spirakis, 2007), $\epsilon\approx 0.5$ for $n$-player polymatrix games, and $\epsilon>0.6$ for general $n$-player games with $n\geq 3$ (Deligkas et al., 2017).   Some special classes of games in which an exact NE can be computed in polynomial time are:
  • Two-player zero-sum games (von Neumann, 1928);
  • Anonymous games in which each player has a fixed small number of strategies and Lipschitz-continuous utility functions (Daskalakis and Papadimitriou, 2007, 2015; Cheng et al., 2017);
  • Graphical games in which the graph (representing the influence between players) is a tree (Kearns et al., 2001) or a general graph with degree 2 (Elkind et al., 2006);
  • Win-lose games with at most 2 winning positions per action of each player (Codenotti et al., 2006);
  • Two-player games in which the payoff matrices have a bounded rank (Lipton et al., 2003; Adsul et al., 2011).
2.3 Aggregative games
An aggregative game is a game in which the utility of each player $i$ depends only on the strategy of $i$, and on some aggregate function of all strategies. Our game corresponds to the special case in which that aggregate function is a sum.
Aggregative games were initially studied for one-dimensional strategy sets (each agent chooses a number) and one-dimensional aggregators. Under certain conditions, such games admit a potential function, [^2] Roughly, a potential function is a function of the strategy profile, whose value increases when any player makes an improving move. A weaker definition only requires the potential to increase when any player makes a best response. and thus always have an NE (Kukushkin, 1994; Dubey et al., 2006; Cornes and Harley, 2012; Jensen, 2018). There are also efficient algorithms for computing an NE when players have discrete strategy sets (Kearns and Mansour, 2002).
For aggregative games with continuous strategy sets, Babichenko (2018) proves that, for every $\epsilon>0$, there exists a best-reply dynamics that reaches an $O(\lambda\epsilon)$-approximate NE in time (number of steps) $O(n\log n/\epsilon^{2})$, where $\lambda$ is the Lipschitz constant of the agents’ utility functions. This result is applicable to a special case of our budget-aggregation game in which there are only $m=2$ issues, as in that case, the strategy of each player is completely determined by the amount he puts on the first issue (the rest of his share automatically goes to the second issue). As far as we know, it is open whether an exact NE for $m=2$ can be computed in polynomial time for this general class of games.
Cummings et al. (2015) study aggregative games with a multi-dimensional aggregator (as in our game), but the players’ strategy sets are finite (unlike our game). For this class, they present an algorithm that computes an NE in time polynomial in the number of players and actions, but exponential in the number of dimensions of the aggregator.
When both players’ strategy sets and the aggregator are multi-dimensional, one can still prove that, under the required assumptions, they have a potential function, and therefore have an NE (Jensen, 2010). However, computing the NE becomes harder. The literature focuses on numeric, distributed algorithms that converge asymptotically to an NE (e.g., Parise et al., 2020; Wang and Nedić, 2024), see Li et al. (2025) for a recent survey. We could not find an algorithm with a polynomial run-time guarantee.
Against this background, our contribution is identifying a natural and practical class of aggregative games, for which an exact NE can be computed efficiently. In the following, different types of aggregative games are covered.
Congestion games
These games are a subclass of aggregative games with multidimensional aggregators, where each coordinate of the aggregator represents the congestion on one of the resources. Usually, the players’ strategy sets are discrete; each player can choose a subset of resources to use (lotteries are not allowed). Equilibria without lotteries are called pure Nash equilibria (PNE).
Even though Nash’s original existence proof is not valid in general for discrete strategy sets, congestion games (CGs) always have a potential function, so every sequence of individual improvements must terminate in a PNE. However, in general, this sequence might be exponentially long. Specifically, computing a PNE in general CGs is PLS-complete, which means that it probably cannot be done by a polynomial-time algorithm (Fabrikant et al., 2004). [^3] PLS is similar to PPAD in that both contain problems in which a solution is guaranteed to exist — both are subsets of TFNP. It is believed that none of them contains the other, and that none of them is contained in P.
Some special classes of CGs in which a PNE can be computed or approximated in polynomial time are:
  • Symmetric network CGs: all players have the same set of possible strategies, and this set is the set of paths between a source node and a target node in a given network (Fabrikant et al., 2004).
  • Nonatomic CGs: CGs in which there is a continuum of players, where each player’s effect on the outcome is negligible (in contrast to our setting, where each player has a substantial effect on the outcome). Finding a PNE in a nonatomic CG can be reduced to solving a convex optimization problem, which can be done in weakly polynomial time. In the special case of a network nonatomic CG with Lipschitz-continuous utility functions, there is an FPTAS (an algorithm computing an $\epsilon$-approximate PNE in time strongly-polynomial in the problem size and $1/\epsilon$). (Fabrikant et al., 2004).
  • CGs in which the strategy space of each player consists of the bases of a matroid over the set of resources (Ackermann et al., 2008).
  • For symmetric (non-network) CGs, an FPTAS based on simple greedy dynamics (Chien and Sinclair, 2011).
  • For general CGs, when the delay functions are polynomials of bounded degree, there are dynamics based on clever sequences of greedy steps that obtain a constant-factor approximate PNE (Caragiannis et al., 2015).
The CG class most closely related to our budget-aggregation game is the class of splittable (atomic) CGs (Orda et al., 2002; Roughgarden and Schoppmann, 2015), in which each agent controls a divisible load and can split it arbitrarily among different paths in a network. The players’ utilities are decreasing functions of the total load allocated to edges that they use: $u_{i}(\delta)=-\sum_{x\in A}\delta_{i,x}\cdot c_{i,x}(\delta_{x})$, where $c_{i,x}$ is the congestion function of edge $x$ for player $i$. An important special case is when the congestion functions are affine, $c_{i,x}(\delta_{x})=a_{i,x}\cdot\delta_{x}+b_{i,x}$. Klimm and Warode (2025) prove that, even with affine congestion functions, computing a PNE is PPAD-complete. Still, polynomial-time algorithms are known for some special cases. If the congestion functions are affine and player-independent, an approximate PNE can be computed by convex programming for general graphs (Cominetti et al., 2009); for specific classes of graphs called ”well-designed”, an exact PNE can be computed by a combinatorial algorithm (Huang, 2013). When strategy sets are singletons — the strategy set of each player is a set of edges, rather than a set of paths from a source to a target — there is a polynomial-time algorithm for affine player-specific congestion functions (Harks and Timmermans, 2022); for convex player-independent congestion functions, there are algorithms that are polynomial when either the number of players or the number of edges is fixed (Bhaskar and Lolakapuri, 2018).
Note that the utility functions are different than our linear utility functions in that the (negative) utility that a player derives from an edge is multiplied by the player’s own contribution to that edge. In other words, the dependence of $u_{i}(\delta)$ on $\delta_{i,x}$ is quadratic, as $\delta_{i,x}$ both multiplies $c_{i,x}$ and contributes to $\delta_{x}$.
Budget-aggregation games with approval preferences
The need to study NEs in budget-aggregation games is also reflected by the fact that many fairness-motivated solution concepts in participatory budgeting and multiwinner voting (see, e.g., Lackner and Skowron, 2023) are related to NEs, especially when agents have approval preferences, i.e., linear utilities with binary valuations, also called dichotomous utilities (Bogomolnaia et al., 2005). For the uncapped setting of budget-aggregation, when projects can receive arbitrary amounts of the budget, Brandl et al. (2022) coined the term “decomposability” for distributions of the overall budget that admit a decomposition into distributions of the individual budgets where each agent spends her budget optimally, i.e., only on approved projects. The standard model in participatory budgeting assumes that projects have fixed costs. For such problems, common rules like the method of equal shares (Peters and Skowron, 2020; Peters et al., 2021) or Phragmén’s rule (Phragmén, 1899; Janson, 2016) rely on the same idea that agents successively allocate their budgets to fund approved projects. Most of these works take a “mechanism design perspective” where agents report their preferences, which are then aggregated into a collective outcome, instead of individual budget distributions. A notable exception is donor coordination (see, e.g., Freeman et al., 2019) where the budget is initially owned by the agents. Brandt et al. (2023, 2025a) proved that, with Leontief utilities, the unique equilibrium distribution is the unique distribution that maximizes the product of agents’ utilities where agents are weighted by their contributions. For the special case of multiwinner approval voting (where all projects have the same costs and correspond to candidates), Haret et al. (2024) recently defined a budgeting game, which is very similar to our budget-aggregation game. In their game, as in ours, each player has a fixed budget, and can choose a distribution of this budget among $m$ issues. But in contrast to our game, their issues are discrete; each issue is active if its total allotment is at least $1$ and inactive otherwise. The utility of a player is the number of active issues that the player approves. They prove that their budgeting game always has an NE, but do not elaborate on its computational aspects. Moreover, returning NEs already ensures some degree of fairness as they satisfy individual fair share (Bogomolnaia et al., 2005; Aziz et al., 2020) for arbitrary preference models (see Proposition 3.4).
2.4 Computing market equilibria
Another equilibrium concept which has been extensively studied from a computational perspective is competitive equilibrium (CE) in markets for private commodities, such as a Fisher market (one seller and many buyers) or the more general exchange market (every agent can both sell and buy; see e.g. Vazirani (2007) for the definitions).
As in the case of NE, the existence of CE in very general settings has been proved using fixed-point theorems (Debreu, 1952; Arrow and Debreu, 1954; McKenzie, 1959); this lead to the development of approximation algorithms without polynomial-time guarantee (Scarf, 1967; Merrill, 1972).
Later, it has been proved that computing an approximate CE is PPAD-hard in general (Papadimitriou, 1994), and even for some special cases such as exchange markets with Leontief utilities (Devanur et al., 2008), exchange markets with separable piecewise-linear concave utilities (Chen et al., 2009), Fisher markets with such utilities, (Chen and Teng, 2009) and exchange markets with negative linear utilities (Chaudhury et al., 2020). Computing an exact CE is FIXP-hard in general (Etessami and Yannakakis, 2010), and even for the special case of Leontief utilities (Garg et al., 2017).
However, CE can be computed in polynomial time for some special cases (some of these cases require either the number of goods $m$ or the number of agents $n$ to be fixed):
  • Exchange markets with Cobb-Douglas utilities (Curtis Eaves, 1985, 2009);
  • Exchange markets with linear utilities when $m$ is fixed (Deng et al., 2002);
  • Fisher markets with Leontief utilities (Codenotti and Varadarajan, 2004);
  • Exchange markets with constant elasticity of substitution (CES), with elasticity at least $1/2$ (Codenotti et al., 2005a);
  • Exchange markets with weak gross substitute utilities (Codenotti et al., 2005b) — a generalization of linear, Cobb-Douglas and CES;
  • Fisher markets with logarithmic utilities when either $n$ or $m$ is fixed (Chen et al., 2004);
  • Exchange markets with piecewise-linear concave utilities when $m$ is fixed (Devanur and Kannan, 2008);
  • Exchange markets with Leontief utilities when $n$ is fixed (Garg et al., 2017);
  • Fisher markets with linear utilities (Devanur et al., 2008; Orlin, 2010);
  • Fisher markets with linear utilities that are negative (Brânzei and Sandomirskiy, 2019) or even a mixture of positive and negative (Garg and McGlaughlin, 2020), when either $m$ or $n$ is fixed;
  • Fisher markets with separable piecewise-linear concave negative functions (Chaudhury et al., 2021).
Most related to our work is an algorithm developed by Jain (2007) for exchange markets with linear utilities; we use some of its theorems in our proofs for Leontief utilities. He first developed a convex program that characterizes all CEs, and then showed how it can be solved efficiently using a variant of the ellipsoid method. Soon afterwards, Ye (2008) showed how to solve this program using interior-point methods, which are more efficient in practice (see also Codenotti et al. (2004) for a survey of similar methods).
3 Model
Let $N$ be a set of $n$ agents and $A$ be a set of $m$ (pure) alternatives. The outcome space $\Delta(A)$ is formed by all distributions $\delta\in[0,1]^{m}$ with $\sum_{x\in A}\delta_{x}=1$. For a distribution $\delta$, we denote its support by $\operatorname{supp}(\delta):=\{x\in A:\delta_{x}>0\}$.
In a budget-aggregation game, the budget (w.l.o.g., normalized to $1$) is equally divided among the agents, and each of them is allowed to distribute their share among the alternatives. The set of strategies available to each player is thus $S_{i}=\{\delta_{i}\in[0,1]^{m}\text{ such that }\sum_{x\in A}\delta_{i,x}=1/n\}$, which is an (m1)(m-1)-dimensional simplex in $\mathbb{R}^{m}$.
Definition 3.1.
A budget-aggregation game is defined as the tuple $(N,A,(u_{i})_{i\in N})$ consisting of
  • a set of agents $N$ with $n=|N|$,
  • a set of alternatives (issues) $A$,
  • and $n$ individual utility functions $u_{i}\colon\Delta(A)\to\mathbb{R}$ for all $i\in N$.
The strategy of agent $i$, denoted $\delta_{i}=(\delta_{i,x})_{x\in A}$, implies that agent $i$ allocates $\delta_{i,x}$ to alternative $x$; these strategies are called individual distributions. We denote the overall distribution by $\delta:=\sum_{i\in N}\delta_{i}$. Note that agents only care about the overall distribution of the budget, which is why the domain of $u_{i}$ is $\Delta(A)$ instead of $S_{1}\times\cdots\times S_{n}$. Furthermore, let $S_{-i}:=S_{1}\times\dots\times S_{i-1}\times S_{i+1}\times\dots\times S_{n}$ be the space of reduced strategy profiles where we exclude agent $i$.
Definition 3.2.
A strategy profile $(\delta^{*}_{i})_{i\in N}$ is called a Nash equilibrium if for all $i\in N$,
$u_{i}(\delta^{*})=\max_{\delta_{i}\in S_{i}}u_{i}\left(\sum_{j\in N\setminus i}\delta^{*}_{j}+\delta_{i}\right)\text{.}$
The sum $\delta^{*}=\sum_{i\in N}\delta^{*}_{i}$ is called an equilibrium distribution.
Interestingly, NEs already guarantee some degree of fairness, as they satisfy a natural fairness notion called individual fair share.
Definition 3.3.
A strategy profile $(\delta_{i})_{i\in N}$ satisfies individual fair share if for all $i\in N$,
$u_{i}\left(\sum_{j\in N}\delta_{j}\right)\geq\max_{\delta^{\prime}_{i}\in S_{i}}\min_{(\delta^{\prime}_{j})_{j\neq i}\in S_{-i}}u_{i}\left(\sum_{j\in N}\delta^{\prime}_{j}\right)\text{.}$
In words, individual fair share ensures that each agent obtains at least as much utility as she could guarantee on her own, by allocating her budget independently of the others.
Proposition 3.4.
Every Nash equilibrium $(\delta^{*}_{i})_{i\in N}$ satisfies individual fair share.
Proof.
For every NE $(\delta^{*}_{i})_{i\in N}$,
$u_{i}(\delta^{*})$ $=\max_{\delta_{i}\in S_{i}}u_{i}\left(\sum_{j\in N\setminus i}\delta^{*}_{j}+\delta_{i}\right)$
$\geq\max_{\delta_{i}\in S_{i}}\min_{(\delta_{j})_{j\neq i}\in S_{-i}}u_{i}\left(\sum_{j\in N\setminus i}\delta_{j}+\delta_{i}\right)\text{.}$
Thus, $(\delta^{*}_{i})_{i\in N}$ satisfies individual fair share. ∎
We are interested in understanding the structure and computing NEs for various utility models. In the following sections, we will consider linear utilities (Section 4), Leontief utilities (Section 5), binary symmetric separable utilities (Section 6), and $\ell_{1}$ disutilities (Section 7).
4 Linear utilities
Linear utilities, also known as additive or von Neumann-Morgenstern utility functions, are the most common extension of valuations for a set of pure alternatives $A$ to preferences over $\Delta(A)$. When agents’ preferences satisfy four seemingly mild conditions, von Neumann and Morgenstern (1944) showed that they can be represented by linear utility functions. Formally, $i$ has linear utilities if for all $\delta\in\Delta(A)$,
$u_{i}(\delta)=\sum_{x\in A}v_{i,x}\delta_{x}.$
with nonnegative valuations $(v_{i,x})_{x\in A}$ for alternatives. [^4] Agents do not have different valuations for individual distributions $\delta_{i,x}$, $\delta_{j,x}$ to the same alternative $x$ as they only care about the total amount $\delta_{x}$ allocated to $x$.
From an economic perspective, alternatives admit an interpretation as substitutes, as agent $i$ is indifferent to how the budget is divided between two alternatives $x,y$ with $v_{i,x}=v_{i,y}$. Due to this property, the set of NEs admits an easily understandable structure. Let $A^{\max}_{i}:=\operatorname*{\arg\max}_{x\in A}v_{i,x}$ be the set of alternatives highest valued by agent $i$.
Proposition 4.1.
For linear utilities, the set of NE consists of all $(\delta_{i})_{i\in N}$ with $\operatorname{supp}(\delta_{i})\subseteq A^{\max}_{i}$ for all $i\in N$.  [^5] If we interpret the overall distribution as a probability distribution, Proposition 4.1 implies that a distribution is in equilibrium if and only if it is the outcome of a random dictatorship, where a player is selected uniformly at random, and then selects one of his best outcomes. Random dictatorship is sometimes used as a benchmark of fairness; see, e.g., Bogomolnaia et al. (2005).
Proof.
By Definition 3.2, $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if $u_{i}(\delta^{*})=\max_{\delta_{i}\in S_{i}}u_{i}(\delta^{*}-\delta^{*}_{i}+\delta_{i})$ for all $i\in N$. By linearity of $u_{i}$, the latter expression is equivalent to $((n-1)/n)\cdot u_{i}(\delta^{*}-\delta^{*}_{i})+(1/n)\max_{\delta_{i}\in S_{i}}u_{i}(n\cdot\delta_{i})$. The term $u_{i}(n\cdot\delta_{i})$ is maximal if and only if $\operatorname{supp}(\delta_{i})\subseteq A^{\max}_{i}$, which proves the statement. ∎
It is straightforward to see that the set of NEs is convex. Its simple structure allows for the computation of NEs with additional properties, e.g., the ones that maximize utilitarian welfare among equilibrium distributions. [^6] Such equilibrium distributions might still be Pareto dominated by other (non-equilibrium) distributions (see, e.g., Brandl et al., 2022).
Furthermore, given a distribution $\delta$, it is easy to check whether it can be decomposed into an NE by solving the following linear feasibility program with variables $d_{i,x}$ (for $i\in N$ and $x\in A$).
$\delta_{x}$ $=\sum_{i\in N}d_{i,x}$ $\text{for all~}x\in A;$
$1/n$ $=\sum_{x\in A^{\max}_{i}}d_{i,x}$ $\text{for all~}i\in N;$
$d_{i,x}$ $\geq 0$ $\text{for all~}i\in N,x\in A.$
5 Leontief utilities
Brandt et al. (2023, 2025a) introduced Leontief preferences for contexts where alternatives share characteristics of complements — meaning that agents only benefit when they receive all of them (in a specific ratio), rather than being able to substitute one for another. For each agent $i$, let $A^{+}_{i}:=\{x\in A\colon v_{i,x}>0\}$ denote the set of alternatives the agent values positively. Leontief preferences are represented by
$u_{i}(\delta)=\min_{x\in A^{+}_{i}}\delta_{x}/v_{i,x}.$
These authors characterize the set of all NEs with the help of the concept of critical alternatives.
Definition 5.1 (Brandt et al. (2025a)).
Given a distribution $\delta\in\Delta(A)$, agent $i$’s set of critical alternatives is defined as
$T_{\delta,i}\coloneqq\arg\min_{x\in A^{+}_{i}}\frac{\delta_{x}}{v_{i,x}}\text{.}$
An alternative is critical in the sense that decreasing its share also decreases the utility of agent $i$.
Lemma 5.2 (Brandt et al. (2023, 2025a), Lemma 1).
In a budget-aggregation game where all agents have Leontief utilities, a profile $(\delta_{i})_{i\in N}$ is an NE if and only if $\operatorname{supp}(\delta_{i})\subseteq T_{\delta,i}$ for all $i\in N$ (where $\delta:=\sum_{i\in N}\delta_{i}$).
Brandt et al. further showed that all NEs result in the same overall distribution (called “equilibrium distribution”), which can be computed by solving the convex program obtained from maximizing the product of agents’ utilities. They leave the question open as to whether the equilibrium distribution can be computed exactly in polynomial time. The remainder of this section is dedicated to answering that question in the affirmative.
Theorem 5.3.
For Leontief utilities, a Nash equilibrium (and the corresponding unique equilibrium distribution) can be computed in time polynomial in the binary encoding length of the input.
[…]
As a first step, we prove an upper bound on the binary encoding length of the equilibrium distribution $\delta^{*}$ and its corresponding utility profile (which we denote by $u^{*}$).
Lemma 5.4.
If the agents’ valuations $v_{i,x}$ are natural numbers, then the equilibrium distribution $\delta^{*}$ and its utility profile $u^{*}$ are rational-valued.   Moreover, their binary encoding length is bounded by a polynomial function of the binary encoding length of $v_{i,x}$ and $n$.
Proof.
For each $i\in N$, let $T_{i}$ be a non-empty set of alternatives. Consider the following linear program (LP), with variables $u_{i}$ (for $i\in N$), $d_{x}$ (for $x\in A$), and $d_{i,x}$ (for $i\in N$ and $x\in T_{i}$):
$d_{x}$ $=u_{i}\cdot v_{i,x}$ $\text{for all~}i\in N,x\in T_{i};$
$d_{x}$ $\geq u_{i}\cdot v_{i,x}$ $\text{for all~}i\in N,x\in A\setminus T_{i};$
$\sum_{x\in T_{i}}d_{i,x}$ $=1/n$ $\text{for all~}i\in N;$
$\sum_{i\in N}d_{i,x}$ $=d_{x}$ $\text{for all~}x\in A;$
$d_{i,x}$ $\geq 0$ $\text{for all~}i\in N,x\in T_{i}\text{.}$
Every solution of this LP (if any) represents a distribution $\delta_{x}=d_{x}$ for all $x\in A$, with a decomposition $\delta_{i,x}=d_{i,x}$ for all $i,x$, such that each agent $i$ contributes only to alternatives in $T_{i}$, and all alternatives in $T_{i}$ are critical for $i$. Variables $u_{i}$ take the values of the corresponding utilities. By Lemma 5.2, such a strategy profile has to coincide with an NE.
An NE $(\delta_{i}^{*})_{i\in N}$ is a solution to the above LP whenever $T_{i}=T_{\delta^{*},i}$ for all $i\in N$. By assumption, the coefficients of this LP are all rational. Therefore, by well-known properties of linear programming, the LP has a rational solution, with binary encoding length bounded by a polynomial function of the representation length of its coefficients.
We can give an explicit bound on the representation length of the equilibrium distribution $\delta^{*}$. For that, we construct an undirected graph where vertices correspond to alternatives, and there is an edge between $x,y\in A$ if and only if there exists an agent $i\in N$ with $x,y\in T_{\delta^{*},i}$. Each component of that graph can be considered separately, as the sum of contributions to that component equals the fraction of agents for whom parts of the component are critical alternatives.
Thus, let $A^{\prime}\subseteq A$ be a subset of $m^{\prime}$ alternatives forming a component and $N^{\prime}\subseteq N$ the subset of agents contributing to alternatives in $A^{\prime}$. Choose some $x_{1}\in A^{\prime}$. If $m^{\prime}=1$ (so $x_{1}$ is the unique alternative in $A^{\prime}$,) then $\delta^{*}_{x_{1}}=N^{\prime}/N$ as $x_{1}$ is the unique critical alternative for all agents in $N^{\prime}$. Otherwise, there needs to be at least one other alternative $x_{2}$ that can be reached in one step from $x_{1}$, i.e., there exists an agent $i_{1}\in N^{\prime}$ such that $\delta^{*}_{x_{1}}/v_{i_{1},x_{1}}=\delta^{*}_{x_{2}}/v_{i_{1},x_{2}}$. Hence, $\delta^{*}_{x_{2}}=(v_{i_{1},x_{2}}/v_{i_{1},x_{1}})\cdot\delta^{*}_{x_{1}}$. Similarly, if $m^{\prime}\geq 3$ then there exists another alternative $x_{3}\in A^{\prime}$ that can be reached in one step from either $x_{1}$ or $x_{2}$. In general, given $k<m^{\prime}$ connected alternatives from $A^{\prime}$, there needs to be another one that can be reached in one step from one of the $k$ alternatives.
This gives a system of linear equations where each $\delta^{*}_{x_{j}}$ can be written in terms of $\delta^{*}_{x_{1}}$. Specifically, $\delta^{*}_{x_{2}}$ only requires the two valuations $v_{i_{1},x_{1}}$ and $v_{i_{1},x_{2}}$, $\delta^{*}_{x_{3}}$ requires at most four valuations ($v_{i_{1},x_{1}},v_{i_{1},x_{2}}$ and $v_{i_{2},x_{2}},v_{i_{2},x_{3}}$ ) , and so on. Considering this “worst case” in terms of representation length together with $\sum_{x\in A^{\prime}}\delta^{*}_{x}=\lvert N^{\prime}\rvert/n$, $\delta^{*}_{x_{1}}$ can be written as a fraction with numerator $\lvert N^{\prime}\rvert/n$ and denominator $1+v_{i_{1},x_{2}}/v_{i_{1},x_{1}}+(v_{i_{1},x_{2}}/v_{i_{1},x_{1}})\cdot(v_{i_{2},x_{3}}/v_{i_{2},x_{2}})+\dots$. This results in a binary encoding length of $\log_{2}((\lvert N^{\prime}\rvert/n)\cdot{v_{\max}}^{m-1})$ for the numerator and $\log_{2}(m\cdot{v_{\max}}^{m-1})$ for the denominator, where we used $m^{\prime}\leq m$, and $v_{i,x}\leq v_{\max}$.
As $x_{1}$ was chosen arbitrarily, and Leontief utilities coincide with distributions to certain alternatives, $(\delta^{*},u^{*})$ can be represented by $n+m$ times the derived length for $\delta^{*}_{x_{1}}$. ∎
Lemma 5.4 cannot be used directly for computing the equilibrium distribution in polynomial time, since the proof requires us to know $T_{\delta^{*},i}$; iterating over all possible $T_{\delta^{*},i}$ would require exponential time. To prove polynomial-time computability, we leverage the following lemma, which is Theorem 13 of Jain (2007):
Lemma 5.5 (Jain (2007)).
Let $S$ be a convex set given by a strong separation oracle, and $\phi>0$ an integer.
There is an oracle-polynomial time and $\phi$-linear time algorithm which does one of the following:
  1. 1.
    Concludes that there is no point in $S$ with binary encoding length at most $\phi$, or
  2. 2.
    produces a point in $S$ with binary encoding length at most $P(n)\cdot\phi$, where $P(n)$ is a polynomial.
We apply Lemma 5.5 as follows. For every positive rational number $z_{0}$, we define a convex set $S(z_{0})\subseteq\mathbb{R}^{n+m}$, where the variables are $u_{i}$ for $i\in N$ and $d_{x}$ for $x\in A$:
$\prod_{i=1}^{n}{u_{i}}^{1/n}$ $\geq z_{0};$
$d_{x}$ $\geq u_{i}\cdot v_{i,x}$  $\text{~for all~}i\in N,x\in A;$
$\sum_{x\in A}d_{x}$ $=1;$
$u_{i}$ $\geq 0$  $\text{~for all~}i\in N;$
$d_{x}$ $\geq 0$  $\text{~for all~}x\in A;$
Intuitively, the set $S(z_{0})$ represents all pairs $(\delta,u)$ such that $\delta$ is a feasible distribution, $u$ is a lower bound on the corresponding utility profile, and the geometric mean of all utilities is at least $z_{0}$; see the proof of Theorem 5.3 below for a formal statement. A strong separation oracle for $S(z_{0})$ is a function that accepts as input a rational vector $\mathbf{y}^{\prime}=(u^{\prime}_{1},\ldots,u^{\prime}_{n},d^{\prime}_{1},\ldots,d^{\prime}_{m})$. It should return one of two outcomes: either an assertion that $\mathbf{y}^{\prime}\in S(z_{0})$, or a hyperplane that separates $\mathbf{y}^{\prime}$ from $S(z_{0})$ (that is, a rational vector $\mathbf{c}$ such that $\mathbf{c}^{\intercal}\mathbf{y}^{\prime}<\mathbf{c}^{\intercal}\mathbf{y}$ for all $\mathbf{y}\in S(z_{0})$).
Lemma 5.6.
For every rational $z_{0}>0$, there is a polynomial-time strong separation oracle for the convex set $S(z_{0})$.
Proof.
Given a rational vector $\mathbf{y}^{\prime}=(u^{\prime}_{1},\ldots,u^{\prime}_{n},d^{\prime}_{1},\ldots,d^{\prime}_{m})$, we first check whether the point satisfies the linear constraints $d^{\prime}_{x}\geq u^{\prime}_{i}\cdot v_{i,x}$, $\sum_{x\in A}d^{\prime}_{x}=1$, $u^{\prime}_{i}\geq 0$ and $d^{\prime}_{x}\geq 0$. If one of these constraints is violated, the constraint itself yields a separating hyperplane. As the number of these constraints is polynomial in $n,m$, all of them can be checked in polynomial time.
It remains to handle the case that all linear constraints are satisfied, whereas the nonlinear constraint is violated, i.e.,
$\prod_{i=1}^{n}u^{\prime}_{i}<{z_{0}}^{n}.$ (1)
As $n\in\mathbb{N}$, the above condition can be checked exactly using arithmetic operations on rational numbers. The binary encoding length of the product is polynomial in the binary encoding length of $u_{i}^{\prime}$, meaning that the representation stays polynomial in the input size.
To construct a separating hyperplane for this case, we use an idea similar to the one in Jain (2007).
First, if $u^{\prime}_{i}=0$ for some $i\in N$, choose $\mathbf{c}$ as the vector with value $1$ for each variable with $u^{\prime}_{i}=0$ and value 0 otherwise. By construction, $\mathbf{c}^{\intercal}\mathbf{y}^{\prime}=0$. For every $\mathbf{y}=(u_{1},\ldots,u_{n},d_{1},\ldots,d_{m})\in S(z_{0})$, $\mathbf{c}^{\intercal}\mathbf{y}>0$ as $\prod_{i=1}^{n}(u_{i})\geq{z_{0}}^{n}>0$.
Otherwise, we can define the vector $\mathbf{c}$ to have the coefficient $(1/n)\cdot(1/u_{i}^{\prime})$ for each variable $u_{i}$, and the coefficient 0 for each variable $d_{x}$. Note that the encoding length of $\mathbf{c}$ is polynomial in the input size. For every vector $\mathbf{y}=(u_{1},\ldots,u_{n},d_{1},\ldots,d_{m})$,
$\mathbf{c}^{\intercal}\mathbf{y}=\frac{1}{n}\cdot\sum_{i=1}^{n}\frac{u_{i}}{u^{\prime}_{i}}.$
Substituting $\mathbf{y}\coloneqq\mathbf{y}^{\prime}$ gives $\mathbf{c}^{\intercal}\mathbf{y}^{\prime}=1$. We now prove that $\mathbf{c}^{\intercal}\mathbf{y}>1$ for every $\mathbf{y}\in S(z_{0})$.
Indeed, $\mathbf{c}^{\intercal}\mathbf{y}$ is the arithmetic mean of the $n$ positive numbers $\frac{u_{i}}{u_{i}^{\prime}}$. By the AM-GM inequality, this sum is at least as large as their geometric mean, that is
$\frac{1}{n}\cdot\sum_{i=1}^{n}\frac{u_{i}}{u^{\prime}_{i}}\;\geq\;\prod_{i=1}^{n}\left(\frac{u_{i}}{u_{i}^{\prime}}\right)^{1/n}\;=\;\frac{(\prod_{i=1}^{n}(u_{i})^{1/n})}{(\prod_{i=1}^{n}(u^{\prime}_{i})^{1/n})}.$
Since $\mathbf{y}=(u_{1},\ldots,u_{n},d_{1},\ldots,d_{m})$ is in $S(z_{0})$, it satisfies the inequality $\prod_{i=1}^{n}u_{i}\geq{z_{0}}^{n}$. Substituting in the right-hand side above gives:
$\frac{1}{n}\cdot\sum_{i=1}^{n}\frac{u_{i}}{u^{\prime}_{i}}\;\geq\;\frac{{z_{0}}}{\prod_{i=1}^{n}(u^{\prime}_{i})^{1/n}},$
which is larger than $1$ by inequality (1). Hence, $\mathbf{c}^{\intercal}\mathbf{y}>1$, so $\mathbf{c}$ indeed defines a separating hyperplane. ∎
Proof of Theorem 5.3.
Lemma 5.6 allows us to apply Lemma 5.5 to $S(z_{0})$. We can now compute the equilibrium distribution by applying binary search to $z_{0}$, as follows.
  1. 1.
    Initialize $L\coloneqq$ the product of agents’ utilities for some arbitrary distribution (e.g., the uniform distribution).
  2. 2.
    Initialize $H\coloneqq$ some upper bound on the product of agents’ utilities, e.g., the one resulting from the (unrealistic) distribution in which each agent $i$ divides the entire budget of $1$ optimally for him (in proportion to $v_{i}$).
  3. 3.
    Let $\phi\coloneqq$ an upper bound on the binary encoding length of $(\delta^{*},u^{*})$, derived in Lemma 5.4.
  4. 4.
    Let $z_{0}\coloneqq(L+H)/2$. Note that both $L$ and $H$ can be encoded in length polynomial in the input size, so the same applies to $z_{0}$.
  5. 5.
    Apply Lemma 5.5 to $S(z_{0})$, using Lemma 5.6 for the strong separation oracle. Consider the two possible outcomes.
    Outcome 1
    (“no point in $S(z_{0})$ with binary encoding length at most $\phi$”): then the vector $(\delta^{*},u^{*})$ is not in $S(z_{0})$. However, the vector $(\delta^{*},u^{*})$ satisfies all linear constraints in the definition of $S(z_{0})$. Hence, it necessarily fails the non-linear constraint. This means that the geometric mean of agents’ utilities in the equilibrium distribution is below $z_{0}$. Set $H\coloneqq z_{0}$ and return to Step 4.
    Outcome 2
    (“a point in $S(z_{0})$ with binary encoding length at most $P(n)\cdot\phi$”): then the $d_{x}$ variables in this point define a distribution $\delta$ by which the utility of each agent $i$ is at least $u_{i}$, and the geometric mean of all $u_{i}$ is at least $z_{0}$; hence, the geometric mean of the agents’ utilities is at least $z_{0}$ too.
    Check whether $\delta$ is the equilibrium distribution (this can be done in polynomial time using Lemma 5.2). If it is, return $\delta$ and finish. Otherwise, set $L\coloneqq z_{0}$ and return to Step 4.
As the binary encoding length of $(\delta^{*},u^{*})$ is at most $\phi$, the binary encoding length of the maximum product of agents’ utilities is at most $\sum_{i}(1/n)\log_{2}(u^{*}_{i})\leq\phi$. Therefore, after at most $\phi$ steps, the binary search is guaranteed to terminate with the equilibrium distribution.
Given the equilibrium distribution, one can easily compute an NE by turning the LP in the proof of Lemma 5.4 into a linear feasibility problem where the variables $u_{i}$ and $d_{x}$ are replaced by their respective values in the equilibrium distribution. ∎
Together with uniqueness of the overall distribution, this implies that the set of NEs is convex.
Moreover, Theorem 5.3 is applicable to other utility models where NEs coincide with the ones for Leontief utilities, e.g., Cobb-Douglas utility functions (Brandt et al., 2025b).
6 Binary symmetric separable utilities
In this section, we investigate NEs for very general utility models when agents have separable preferences. Separability requires that agents receive utility independently from each alternative. More formally, fixing the outcome on a subset of all alternatives has no influence on an agent’s preferences over how to allocate the remaining budget among the other alternatives. The interested reader is advised to consult Debreu (1959); Blackorby et al. (1978) for an axiomatic characterization of separable preferences. Here, we say that preferences are separable if they can be represented by a utility function of the form
$u_{i}(\delta)=\sum_{x\in A}g_{i,x}(\delta_{x})$
where $g_{i,x}\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ are continuous and weakly increasing.
Linear utilities represent the most common class of separable preferences. Contrarily, Leontief preferences are not separable as the minimum is taken over all alternatives with positive valuations.
We consider the class of binary symmetric separable preferences where each agent $i$ has a non-empty set of approved alternatives $A_{i}\subseteq A$ but does not further distinguish between the alternatives. [^7] This is what we mean by the term “symmetric” in our context. Furthermore, her utility is assumed to strictly increase when the budget on some $x\in A_{i}$ increases.
In terms of utility functions, such preferences are represented by
$u_{i}(\delta)=\sum_{x\in A_{i}}g_{i}(\delta_{x})$
where $g_{i}\colon[0,1]\to\mathbb{R}$ are continuous and strictly increasing functions.
For example, setting $g_{i}(\delta_{x})=\delta_{x}$ for all $i\in N$ corresponds to linear preferences with binary valuations $v_{i,x}\in\{0,1\}$ (see, e.g., Bogomolnaia et al., 2005). Choosing $g_{i}(\delta_{x})=\ln(\delta_{x})$ leads to Cobb-Douglas utilities with binary weights. However, note that our model allows the $g_{i}$’s to differ among agents.   The following example shows that for such general preferences, NEs do not always exist.
Example 6.1.
Consider an instance with $n=2$ and $A=\{a,b\}$, with $g_{1}(\cdot)=\sqrt{\cdot}$, i.e., $u_{1}(\delta)=\sqrt{\delta_{a}}+\sqrt{\delta_{b}}$ and $g_{2}(\cdot)=\cdot^{2}$, i.e, $u_{2}(\delta)=\delta_{a}^{2}+\delta_{b}^{2}$.
If $\delta_{a}>\delta_{b}$, then $\delta_{a}>1/2$, hence both agents contribute to $a$. By strict concavity of $g_{1}$, Agent $1$ has an incentive to move some budget from $a$ to $b$. The case $\delta_{a}<\delta_{b}$ can be treated analogously.
If $\delta_{a}=\delta_{b}$, then by strict convexity of $g_{2}$, Agent $2$ has an incentive to move her entire budget to an alternative $x\in\{a,b\}$ with $\delta_{1,x}\geq 1/4$, resulting in $\delta_{a}\neq\delta_{b}$. Thus, there cannot exist an NE.
Intuitively, Agent $1$ wants the entire budget to be distributed as equally as possible, whereas Agent $2$ prefers to focus on one alternative.   Therefore, we treat separately the two cases where all $g_{i}$’s are strictly concave (Section 6.1) and strictly convex (Section 6.2).
6.1 Strictly concave utilities
In this section, all $g_{i}$’s are assumed to be strictly concave, i.e., $g_{i}(\lambda\delta+(1-\lambda)\delta^{\prime})>\lambda g_{i}(\delta)+(1-\lambda)g_{i}(\delta^{\prime})$ for all different $\delta,\delta^{\prime}\in\Delta(A)$ and $\lambda\in(0,1)$.
Theorem 6.2 (Brandt et al. (2025a), Proposition 9).
For binary symmetric separable and strictly concave utilities, the set of NEs coincides with that for Leontief utilities with the same valuations ($v_{i,x}\in\{0,1\}$).
For binary valuations, Brandt et al. (2025a) stated two different possibilities to compute an overall equilibrium distribution via linear programming. This distribution can be decomposed into an NE using the same linear program described at the end of Theorem 5.3.
Corollary 6.3.
For binary symmetric separable and strictly concave utilities, an NE can be computed via linear programming.
Furthermore, the concept of critical alternatives can readily be transferred in order to check whether a given strategy profile constitutes an NE.
6.2 Strictly convex utilities
In this section, all $g_{i}$’s are assumed to be strictly convex, i.e., $\lambda g_{i}(\delta)+(1-\lambda)g_{i}(\delta^{\prime})>g_{i}(\lambda\delta+(1-\lambda)\delta^{\prime})$ for all different $\delta,\delta^{\prime}\in\Delta(A)$ and $\lambda\in(0,1)$.
Note that in this case, each utility function $u_{i}(\delta)$, seen as an $m$-dimensional real-valued function, is convex, as it is a sum of strictly convex functions. Furthermore, it is strictly convex if the domain is restricted to the $\lvert A_{i}\rvert-$dimensional subspace spanned by agent $i$’s approved alternatives.
Let $M_{\delta,i}\coloneqq\arg\max_{x\in A_{i}}\delta_{x}$ be the set of alternatives with maximum share under the distribution $\delta$ that are approved by agent $i$.
Lemma 6.4.
For binary symmetric separable and strictly convex utilities, the set of Nash equilibria consists of all $(\delta^{*}_{i})_{i\in N}$ with $\operatorname{supp}(\delta^{*}_{i})=M_{\delta^{*},i}$ and $\lvert\operatorname{supp}(\delta^{*}_{i})\rvert=\lvert M_{\delta^{*},i}\rvert=1$ for all $i\in N$.
Proof.
The distribution $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if every agent $i$’s distribution $\delta^{*}_{i}$ is a best response and therefore a solution of the optimization problem
$\max_{\delta_{i}\in S_{i}}u_{i}(\delta^{*}-\delta^{*}_{i}+\delta_{i})\text{.}$
The objective function is strictly convex on this domain, and the feasible region is the ($\lvert A_{i}\rvert-1$)-dimensional simplex spanned by those strategies where agent $i$ places her entire budget on a single approved alternative. So every maximizer is a vertex of this simplex, which proves that $\lvert\operatorname{supp}(\delta_{i})\rvert=1$. By adding $1/n$ to an alternative $x\in A_{i}$ in $\delta^{*}-\delta^{*}_{i}$, agent $i$’s utility increases by $g_{i}(\delta^{*}_{x}-\delta^{*}_{i,x}+1/n)-g_{i}(\delta^{*}_{x}-\delta^{*}_{i,x})$. This increase is maximal if and only if $x$ is an alternative with maximal share under the distribution $\delta^{*}-\delta^{*}_{i}$, i.e., $x\in M_{\delta^{*}-\delta^{*}_{i},i}$ since $g_{i}$ is strictly convex. This proves that $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if every agent plays a pure strategy, i.e., $\lvert\operatorname{supp}(\delta^{*}_{i})\rvert=1$, and $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{\delta^{*}-\delta^{*}_{i},i}$ for all $i\in N$. Thus, $\lvert M_{\delta^{*},i}\rvert=1$ and $\operatorname{supp}(\delta^{*}_{i})=M_{\delta^{*},i}$ for all $i\in N$.
For the other direction, it remains to show that if $\operatorname{supp}(\delta_{i})=M_{\delta,i}$ and $\lvert M_{\delta,i}\rvert=1$ for all $i\in N$, it holds that $\operatorname{supp}(\delta_{i})\subseteq M_{\delta-\delta_{i},i}$ for all $i\in N$. Therefore, assume $x\in\operatorname{supp}(\delta_{i})$ and $y\in A\backslash\{x\}$ for agent $i$. Then, since all agents play pure strategies, all $\delta_{x}$ are some multiple of $1/n$ and $\delta_{x}\geq\delta_{y}+1/n$. Thus,
$\delta_{y}-\delta_{i,y}=\delta_{y}\leq\delta_{x}-1/n=\delta_{x}-\delta_{i,x}$
and $\operatorname{supp}(\delta_{i})\subseteq M_{\delta-\delta_{i},i}$. ∎
Corollary 6.5.
For a Nash equilibrium $(\delta^{*}_{i})_{i\in N}$ and two alternatives $x,y\in A$, $\delta^{*}_{x}=\delta^{*}_{y}$ implies $\lvert A_{i}\cap\{x,y\}\rvert=1$ for all $i\in N$ with $\operatorname{supp}(\delta^{*}_{i})\in\{x,y\}$.
Lemma 6.4 implies that checking whether a given distribution $\delta$ admits an NE $(\delta_{i})_{i\in N}$ can be done in polynomial time. In a first step, verify that $\lvert M_{\delta,i}\rvert=1$ for all $i\in N$. Second, check that the decomposition $(\delta_{i})_{i\in N}$ with $\delta_{i,x}=1/n$ for $x\in M_{\delta,i}$ (and $\delta_{i,x}=0$ otherwise) is a valid decomposition of $\delta$.
Theorem 6.6.
For binary symmetric separable and strictly convex utilities, a Nash equilibrium always exists and can be found in polynomial time.
Proof.
We use the following greedy approach to construct an NE $(\delta_{i}^{*})_{i\in N}$.  [^8] This algorithm is equivalent to an algorithm called “maximum payment rule” (Aziz et al., 2025), which was recently presented as a rule for approximately-fair budget distribution for agents with binary linear utilities. Let $N^{\prime}\subseteq N$ denote the set of agents who have not yet allocated their budgets. In each step, select an alternative $x^{*}\in\arg\max_{x\in A}\sum_{i\in N^{\prime}:x\in A_{i}}1/n$, i.e., $x^{*}$ maximizes the aggregated share of the remaining agents who approve it. Each of these agents then allocates her entire budget to $x^{*}$, and is excluded from $N^{\prime}$. The procedure is repeated until no agents remain.
This process runs in polynomial time in $n$ and $m$ as it terminates after at most $n$ steps with a complete strategy profile $(\delta^{*}_{i})$ where in each step, a suitable alternative $x^{*}$ can be found by counting approvals of the remaining agents in $N^{\prime}$. By construction, $\lvert\operatorname{supp}{\delta^{*}_{i}}\rvert=1$ for all $i\in N$. Furthermore, $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{\delta^{*},i}$. Otherwise, some agent $i$ approves an alternative $x$ with $\delta^{*}_{x}>\delta^{*}_{\operatorname{supp}(\delta^{*}_{i})}$ and would have contributed to $x$ instead under the described greedy approach.
Finally, if $\lvert M_{\delta^{*},i}\rvert>1$ for some agent $i$, then there exist alternatives $x,y\in A_{i}$ with $\delta^{*}_{x}=\delta^{*}_{y}$ and w.l.o.g. $\operatorname{supp}(\delta^{*}_{i})=x$, contradicting Corollary 6.5.
Thus, $(\delta^{*}_{i})_{i\in N}$ constitutes an NE by Lemma 6.4. ∎
The following example shows that the set of NEs is not convex in general.
Example 6.7.
Consider a profile with a single agent $i$ and two alternatives $x$ and $y$ that are both approved by the agent, i.e., $A_{i}=A$. Using Lemma 6.4, it is easy to see that there exist exactly two equilibria, which are given by the two distributions that allocate the entire budget to one of the alternatives.
In fact, the number of NE is finite and can be bounded explicitly.
Lemma 6.8.
The size of the set of Nash equilibria is bounded by $\min\{m!,\prod_{i\in N}\lvert A_{i}\rvert\}$
Proof.
First, there are $m!$ possibilities to strictly order the set of alternatives by the amount of budget they receive. For each ordering, there can exist at most one NE for which the total distribution follows this ordering from highest to lowest share placed on the alternatives. Ties can be broken arbitrarily due to Corollary 6.5. By Lemma 6.4, the ordering uniquely determines the sets $M_{\delta,i}$ and thus, the possible corresponding NE.
Second, Lemma 6.4 shows that in any NE, every agent always plays a pure strategy. This implies that each agent $i$ has at most $\lvert A_{i}\rvert$ different best responses. Therefore, there are $\prod_{i\in N}\lvert A_{i}\rvert$ possible combinations of best responses, i.e., potential NEs. ∎
Both of these upper bounds are sharp for certain profiles.
Example 6.9.
First, consider a profile with $m$ alternatives and $n:=2^{m}-1$ agents, one agent for every possible set of approved alternatives. In particular, each alternative is approved by half of the agents. Then, for every possible strict ordering of alternatives $x_{1}\succ\dots\succ x_{n}$, there exists an equilibrium with shares following this ordering: alternative $x_{\ell}$ receives the budgets of all agents $i$ with $x_{\ell}\in A_{i}$ and $x_{k}\not\in A_{i}$ for all $k<\ell$, i.e., $2^{m-1}/2^{\ell-1}=2^{m-\ell}$ agents. Thus, the allocated shares in decreasing order are given by $2^{m-1}/n,2^{m-2}/n,...,1/n$.
Second, for every set of positive integers $a_{1},\dots,a_{n}$, consider a profile with $n$ agents and $m:=a_{1}+\cdots+a_{n}$ alternatives, such that each agent $i$ approves a set $A_{i}$ of $a_{i}$ alternatives, and all $A_{i}$’s are pairwise disjoint. Then, every possible combination of strategies $\delta_{i}$ with $\lvert\operatorname{supp}(\delta_{i})\rvert=1$ and $\operatorname{supp}(\delta_{i})\in A_{i}$ results in an equilibrium, giving $\prod_{i=1}^{n}\lvert A_{i}\rvert$ equilibria in total.
7 $\ell_{1}$ disutilities
A widely studied utility model in the context of budget-aggregation is based on $\ell_{1}$ preferences, also referred to as $\ell_{1}$ disutilities (Lindner, 2011; Goel et al., 2019; Freeman et al., 2019). The existing literature mostly focuses on the axiomatic investigation of mechanisms (Brandt et al., 2024; Elkind et al., 2026), in particular, moving phantom mechanisms (Freeman et al., 2021; de Berg et al., 2024).
For $\ell_{1}$ preferences, each agent is assumed to have a peak allocation — a preferred distribution of the budget — and her utility is determined by the negative $\ell_{1}$ distance between any given allocation and her peak. Formally, let $p_{i}=(p_{i,x})_{x\in A}\in\Delta(A)$ denote the peak of agent $i$, then her utility for a given allocation $\delta\in\Delta(A)$ is defined as
$u_{i}(\delta)=-\lVert\delta-p_{i}\rVert_{1}.$
Alternatively, the utility can be interpreted in terms of overlap utilities, as shown by Goel et al. (2019). In the overlap utility model, the utility of agent $i$ is defined as the sum of the overlaps between the allocation $\delta$ and the agent’s peak $p_{i}$:
$u_{i}(\delta)=\sum_{x\in A}\min(\delta_{x},p_{i,x}).$
Goel et al. (2019) demonstrate that $\ell_{1}$ disutilities are equivalent (up to an affine transformation) to overlap utilities, making both models interchangeable in our budget-aggregation context. Note that these preferences are separable, but not binary symmetric separable.
These utilities capture the behavior of agents who derive value from receiving resources up to a certain satiation point, beyond which additional resources provide no further benefit. Specifically, the utility is linear in the allocation up to the agent’s peak, and constant thereafter — reflecting satiation once the desired threshold is met.
We say an alternative $x$ is oversupplied for agent $i$ if $\delta_{x}>p_{i,x}$. Additionally, let $M_{i}^{>}(\delta)$ be the set of alternatives that are oversupplied for agent $i$ given an allocation $\delta$. Similarly, an alternative $x$ is undersupplied for agent $i$ if $\delta_{x}<p_{i,x}$ and $M_{i}^{<}(\delta)$ is the set of undersupplied alternatives. Define $M_{i}^{\leq}(\delta^{*})$ accordingly. We show that a strategy profile constitutes an NE when no agent contributes to one of her oversupplied alternatives.
Lemma 7.1.
For $\ell_{1}$ preferences, a strategy profile $(\delta^{*}_{i})_{i\in N}$ is a Nash equilibrium if and only if $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{i}^{\leq}(\delta^{*})$ for all $i\in N$.
Proof.
$\Rightarrow$: Assume that there is some agent $i$ with $\delta^{*}_{i,x}>0$ for some alternative $x$ which is oversupplied, i.e. $\delta^{*}_{x}>p_{i,x}$. Redistributing $\delta^{*}_{i,x}$ to some undersupplied alternative, say $x^{\prime}$, strictly increases the utility of agent $i$, showing that the strategy profile $(\delta^{*}_{i})_{i\in N}$ is not an NE.
$\Leftarrow$: Suppose $(\delta^{*}_{i})_{i\in N}$ satisfies $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{i}^{\leq}(\delta^{*})$ for all $i\in N$. Since the marginal utility decrease in $x$ equals the marginal utility increase in $x^{\prime}$ for all $x,x^{\prime}\in M_{i}^{\leq}(\delta^{*})$, agent $i$ cannot strictly increase her utility by redistributing her budget $\delta_{i}$. Therefore, $u_{i}(\delta^{*})=\max_{\delta_{i}^{\prime}\in S_{i}}u_{i}(\delta^{*}-\delta^{*}_{i}+\delta_{i}^{\prime})$, establishing that $(\delta^{*}_{i})_{i\in N}$ is an NE. ∎
The existence of an NE for $\ell_{1}$ disutilities follows from the more general existence result by Debreu (1952) for quasi-concave utilities. Given our above characterization, it is not surprising that NEs do not yield a unique overall distribution in the case of $\ell_{1}$ preferences, since the characterization depends solely on the support of the agents’ individual distributions.
Example 7.2.
Let $n=3$ and consider the following profile
$P=\begin{pmatrix}3/10&3/10&0&4/10\\ 2/10&2/10&6/10&0\\ 2/10&2/10&6/10&0\\ \end{pmatrix}$,
where each row denotes the peak of an agent. Since this instance has three agents, everyone has a budget of $1/3$. Both $\delta=(3/10,2/10,14/30,1/30)$ and $\delta^{\prime}=(2/10,3/10,14/30,1/30)$ admit decompositions into strategies that satisfy Lemma 7.1, i.e. no agent contributes to oversupplied alternatives:
$\left(\delta_{i}\right)_{i\in N}=\begin{pmatrix}3/10&0&0&1/30\\ 0&2/10&4/30&0\\ 0&0&10/30&0\\ \end{pmatrix}$
$\left(\delta^{\prime}_{i}\right)_{i\in N}=\begin{pmatrix}0&3/10&0&1/30\\ 2/10&0&4/30&0\\ 0&0&10/30&0\\ \end{pmatrix}$
This verifies that both agent profiles $(\delta_{i})_{i\in N}$ and $(\delta_{i}^{\prime})_{i\in N}$ are NEs for the given set of peaks $P$.
With the above example, we can show the following.
Proposition 7.3.
For $\ell_{1}$ preferences, the set of equilibrium distributions is not convex.
Proof.
For the proof, consider again Example 7.2 and a convex combination of the two overall distributions $\delta$ and $\delta^{\prime}$. More specifically, let $\delta^{\prime\prime}:=\frac{1}{2}\delta+\frac{1}{2}\delta^{\prime}$. We get $\delta^{\prime\prime}=(1/4,1/4,14/30,1/30)$. Since Agents $2$ and $3$ only support the first and second alternative up to $2/10$, every strategy profile that is in equilibrium must have $\delta_{i,x}=0$ for $i\in\{2,3\}$ and $x\in\{1,2\}$. Since the total budget of the first agent is only $1/3$, she cannot cover both alternatives simultaneously. Consequently, $\delta^{\prime\prime}$ cannot be decomposed into a strategy profile that is also an NE. This establishes that the set of equilibrium distributions is not convex. ∎
Although the set of NEs may be non-convex — demonstrating potential structural complexity — there nonetheless exists a polynomial-time algorithm that computes an NE.
Theorem 7.4.
For $\ell_{1}$ preferences, a Nash equilibrium can be computed in polynomial time.
Proof.
We present a polynomial-time algorithm. Let each agent $i$ have a budget, denoted by $b_{i}$. Initially, this budget is set to $b_{i}=1/n$ for all $i\in N$. Iterate over all alternatives $x\in A$ in an arbitrary order. For a given alternative $x$, order all agents in a descending order of their peaks for this alternative, i.e., $p_{i_{1},x}\geq p_{i_{2},x}\geq\dots\geq p_{i_{n},x}$. When it is agent $i$’s turn, she allocates $\min(b_{i},\max(0,p_{i,x}-\delta_{x}))$ on alternative $x$. Afterwards, $b_{i}$ is decreased and $\delta_{x}$ is increased by this amount. Once agent $i$ has spent her entire share ($b_{i}=0$), she is deleted from the set of considered agents. The ordering of the agents ensures that $x$ does not become oversupplied to an agent who already contributed to $x$ later on, i.e., no agent contributes to an oversupplied alternative. When no agent wants to contribute to alternative $x$, we go to the next alternative $x^{\prime}$ and reorder the remaining agents according to their peaks ($p_{i_{1},x^{\prime}}\geq p_{i_{2},x^{\prime}}\dots\geq p_{i_{n},x^{\prime}}$). Then, we let them again sequentially contribute to this alternative.
We prove that every agent spends her entire budget after iterating over all alternatives. Assume for contradiction that there is some agent who has still some budget left after she had the possibility to contribute to every alternative. This would mean that $p_{i,x}\leq\delta_{x}$ for all $x\in A$ in the final partial distribution $\delta$, which is impossible due to the fact that $p_{i}\in\Delta(A)$ but $\sum_{x\in A}\delta_{x}<1$. All in all, the suggested construction returns a strategy profile that constitutes an NE.
Iterating over all $m$ alternatives and, for every alternative, sorting at most $n$ agents according to their peaks leads to a run-time complexity of $O(mn\log n)$. ∎
This construction provides an efficient algorithm for obtaining an NE given a set of optimal distributions of the agents. Additionally, we can also show whether a given overall distribution can be represented as individual distributions that constitute an NE.
Theorem 7.5.
For $\ell_{1}$ preferences, one can check in polynomial time whether an overall distribution can be decomposed into a Nash equilibrium.
Proof.
Let $\delta$ be the overall distribution. We now consider the following linear constraints to find individual distributions constituting an NE where $d_{i,x}$ are the variables for all $i\in N$ and $x\in A$:
$\sum_{x\in A}d_{i,x}\leq\frac{1}{n}\text{ for all }i\in N,$
$\sum_{i\in N}d_{i,x}=\delta_{x}\text{ for all }x\in A,$
$d_{i,x}(p_{i,x}-\delta_{x})\geq 0\text{ for all }i\in N\text{ and }x\in A,$
$d_{i,x}\geq 0\text{ for all }i\in N\text{ and }x\in A.$
Let $d_{i,x}$ be the individual distribution of agent $i$ for alternative $x$, i.e., $d_{i,x}=\delta_{i,x}$ for all $i\in N$ and $x\in A$. The third constraint requires that an agent never contributes to an oversupplied alternative, which is a necessary and sufficient condition for a set of strategy profiles to form an NE (see Lemma 7.1). Furthermore, the second constraint ensures that individual strategies $(\delta_{i})_{i\in N}$ form the overall distribution $\delta$. Therefore, if these constraints are satisfied, the individual strategies add up to the overall distribution, which is in turn an equilibrium distribution due to the third constraint. ∎
8 Weighted agents
In practice, one might want to assign different shares to the agents, e.g., when they represent groups of citizens of different sizes, or citizens in different tax brackets.
We define a weighted budget-aggregation game as a tuple $(N,A,\allowbreak(u_{i})_{i\in N},(w_{i})_{i\in N})$, where $N$, AA and $(u_{i})_{i\in N}$ are as in Definition 3.1, and $w_{i}$ is the weight of agent $i$, with $\sum_{i\in N}w_{i}=1$. The strategy set of each agent $i$ is given by
$S_{i}(w_{i}):=\{\delta_{i}\in[0,1]^{m}\text{ such that }\sum_{x\in A}\delta_{i,x}=w_{i}\}\text{.}$
The definition of an NE transfers directly from Definition 3.2, except that $S_{i}$ depends on $w_{i}$ as stated above.
For some classes of utility functions, the weighted setting can be reduced to the unweighted setting by agent cloning if all weights are rational numbers, resulting in a pseudo-polynomial time algorithm. We formalize the agent cloning process below.
Definition 8.1.
Given a weighted budget-aggregation game $(N,A,\allowbreak(u_{i})_{i\in N},(w_{i})_{i\in N})$ where all $w_{i}$ are rational numbers, let $D$ be their least common denominator, such that $w_{i}=q_{i}/D$ for all $i\in N$, and all $q_{i}$ are integers. The corresponding cloned game is an (unweighted) budget-aggregation game in which the set of agents $N^{c}$ contains $q_{i}$ clones of each agent $i$ (with the same utility function $u_{i}$), so that $|N^{c}|=\sum_{i\in N}q_{i}=D$.
Lemma 8.2.
Let $(N,A,(u_{i})_{i\in N},(w_{i})_{i\in N})$ be a weighted budget-aggregation game in which all $w_{i}$ are rational numbers, and all $u_{i}$ are either linear, Leontief, binary symmetric separable and strictly concave/convex, or $\ell_{1}$. Let $(\delta_{i^{c}})_{i^{c}\in N^{c}}$ be a Nash equilibrium of the corresponding cloned game. Define a strategy profile $(\delta_{i})_{i\in N}$ such that $\delta_{i}$ is the sum of $\delta_{i^{c}}$ for all $i^{c}$ which are clones of $i$. Then $(\delta_{i})_{i\in N}$ is a Nash equilibrium of the weighted game.
Proof.
Suppose that $(\delta_{i})_{i\in N}$ is not an NE of the weighted game. For linear, Leontief, and $l_{1}$ preferences, there is a structural lemma that implies that, under $\delta$, some agent $i$ contributes a positive amount to some project $y$ in a set of “forbidden” projects that depends only on $u_{i}$ and $\delta$. Hence, under $\delta$, some clone $i_{c}$ of $i$ must contribute a positive amount to the same project $y$. As $i_{c}$ and $i$ have the same utility function, the set of “forbidden” projects under $\delta$ is the same for both of them; as $i_{c}$ contributes to a “forbidden” project, $\delta$ does not correspond to an NE of the cloned game.
Specifically:
  • For linear utilities, by Proposition 4.1, some agent $i$ contributes under $\delta$ to some non-highest-valued project $y\in A\setminus A^{\max}_{i}$;
  • For Leontief utilities, by Lemma 5.2, some agent $i$ contributes under $\delta$ to some non-critical project $y\in A\setminus T_{\delta,i}$;
  • For $\ell_{1}$ disutilities, by Lemma 7.1, some agent $i$ contributes under $\delta$ to some over-supplied project $y\in M^{>}_{i}(\delta)$.
It remains to prove the lemma for binary symmetric separable and strictly concave/convex preferences.
Theorem 6.2 remains true for the weighted game (Brandt et al., 2025a). Thus, some agent in the cloned game has to contribute to a non-critical alternative under $(\delta_{i})_{i\in N}$ in the cloned game and has a beneficial deviation, showing that the statement also holds for binary symmetric separable and strictly concave preferences.
Finally, if $(\delta_{i})_{i\in N}$ is no NE of the weighted game for binary symmetric separable and strictly convex preferences, this means that some agent $i$ contributes either to an alternative outside of $M_{\delta,i}$ or to two alternatives $x,y\in M_{\delta,i}$. For the first case, this implies that there exists a clone of agent $i^{c}$ that contributes to an alternative outside of $M_{\delta,i^{c}}$ in the cloned game. Corollary 6.5 proves that also for the second case, $(\delta_{i})_{i\in N}$ cannot refer to an NE of the cloned game.∎
Theorem 8.3.
A Nash equilibrium can be computed in pseudo-polynomial time for weighted budget-aggregation games in which all utilities are either linear, Leontief, binary symmetric separable and strictly concave/convex, or $\ell_{1}$.
Proof.
Given a weighted budget-aggregation game with weights $w_{i}=q_{i}/D$, construct the corresponding cloned game, which is an unweighted game. This game has $D$ agents; hence, an NE can be computed in time polynomial in $m$ and $D$. By Lemma 8.2, this NE corresponds to an NE for the weighted game; the run-time is pseudo-polynomial in the parameters of that game. ∎
We believe Lemma 8.2 can be extended to other classes of utilities. It would be interesting to characterize the set of utility functions for which Lemma 8.2 remains true. Moreover, it is open whether an NE in a weighted budget-aggregation game can be computed in time polynomial in $n$ and $m$ (rather than pseudo-polynomial), e.g., for Leontief preferences.
9 Conclusion
In this paper, we introduced budget-aggregation games to formalize a game-theoretic distributed approach for dividing a budget among a set of public projects and investigated NEs of the resulting games. Although finding NEs is known to be hard for general classes of utility models, our work shows that in practice, they can be computed efficiently. In detail, we proved that for the most common preference models of linear, Leontief, and $\ell_{1}$ preferences as well as for classes of binary symmetric and separable preferences, NEs can be computed in polynomial time.
It would be interesting to see whether there exists a class of utility functions containing all of the previously considered preference models that still allow for an efficient computation of NEs. As a first step, one might want to extend binary symmetric separable utilities to allow for valuations other than 0 or $1$, or consider preferences based on $\ell_{p}$ metrics other than $\ell_{1}$.
Other possible directions for future work include the selection of specific equilibria, e.g., by incorporating additional axioms like Pareto efficiency or investigating limits of equilibrium dynamics, and the investigation of the price of anarchy.
Acknowledgments
We thank Yakov Babichenko, Felix Brandt, Warut Suksompong, and the anonymous reviewers of AAMAS 2026 for helpful comments. Patrick is supported by the Deutsche Forschungsgemeinschaft under grant BR 2312/14-1. Erel is supported by the Israel Science Foundation grant no. 1092/24.